OpenStudy (anonymous):

I am soooo very desperate I need help with 3 problems I've been stuck on for an hour please help and don't guess I need to make sure they are 100% right ??!?

OpenStudy (anonymous):

which ones

OpenStudy (anonymous):

What are the explicit equation and domain for a geometric sequence with a first term of 4 and a second term of -12?

OpenStudy (anonymous):

sorry I can't help

OpenStudy (perl):

how do you get from the first term to the second term?

OpenStudy (anonymous):

2. What are the explicit equation and domain for an arithmetic sequence with a first term of 5 and a second term of 2?

OpenStudy (anonymous):

@perl I'm not sure I just really need help and to get answers asap please

OpenStudy (anonymous):

And 3. What is the sum of the geometric sequence -3,24,-144 if there are 6 terms? I think this one is 26,661 but I'm not sure

OpenStudy (anonymous):

@dddd47906 could you help me on the last one maybe ?

OpenStudy (perl):

for geometric sequence , the explicit equation is $$ \LARGE { a_n= a_1\cdot r^{n-1} }$$

OpenStudy (anonymous):

@perl okay and than ?(:

OpenStudy (perl):

r is the ratio of consecutive terms r = -12/4 = -3

OpenStudy (anonymous):

I am super lost ;( what you by an chance show me the steps and answer all In one for each of the question pretty please you would seriously be a life savior

OpenStudy (anonymous):

@perl

OpenStudy (perl):

I'm not allowed to give the answer, but I can guide you.

OpenStudy (anonymous):

Um, I haven't learned that... You see, I'm still in 6th grade.

OpenStudy (anonymous):

@perl I know your not to just hand me the answer but could you show me the steps to get the answer please